A 2.0 kg block is sliding across a table with an initial speed of 0.5 m/s. The table imparts a 20.0 N force due to friction on the block. How much work in Joules will the table do on the block?
Answer #1

at Im=2.0kg mass of the block. u=a.5m/s) ng » Initial speed. v=omis > Fimal speed. work done by normal and ng will be Zero, bImrz f 1 / 2 mu? 2 > Wf t Wrt Wmg → Wp +0+0=0- (0 - 1x2x60.50) 3 >> -- 0.25 w 0.25 I